# Learning to Route LLMs with Confidence Tokens

### v2

@@ 5.1 Routing Performance (RQ1)
Figure 3: ROC curves of Llama3-8B-Instruct on MMLU (a) and OpenbookQA (b); and Mistral-7B-Instruct on MMLU (c) and OpenbookQA (d), for the rejection learning task. The true positive rate (recall) indicates the proportion of samples where “none of the above” is the ground truth that are correctly predicted as such. The false positive rate indicates the proportion of original samples that have been falsely predicted for rejection. Across all settings, Self-REF outperforms other baselines.
@@ 5.2 Rejection Learning Performance (RQ2)
In this section, we analyze the utility of confidence scores from Self-REF for the purpose of learning to abstain from answering. Without incorporating additional knowledge of the rejection task during training, Self-REF and baselines rely solely on thresholding the confidence scores to determine whether to reject a prediction. In safety-critical applications where one may hope to detect cases should bed abstained from answering (i.e., high recall on the reject option) while retaining performance on the task at hand (i.e., low false positive rate), it is useful to examine the ROC curve, which makes this trade-off. Across all thresholds, we observe that on both MMLU and OpenbookQA dataset, Self-REF outperforms all baselines on the rejection learning task (Figure 3). Because the confidence tokens are trained and conditioned on the correctness of the predicted answers, Self-REF has a higher likelihood of being aware of the correctness of its generated responses using these tokens. In this manner, Self-REF may better distinguish when there is no correct answer to choose, resulting in better rejection learning capabilities.
@@ 5.3 Calibration with Utility (RQ3)
Here, we analyze the correlation between calibration, or how well predicted probabilities align with actual probabilities, and utility for downstream routing. Following the settings from Tian et al. [48], Błasiok et al. [4], three calibration metrics are used to assess the calibration: (1) expected calibration error, ECE [14]; (2) Brier Score, BS [5]; and (3) the cross-entropy score, CE. Alongside these calibration metrics, we also show the least rates of routing required for each technique to attain comparable performance to Llama3-70B-Instruct. As shown in Table 2, although Self-REF achieves the best or second-best calibration scores, better-calibrated methods do not guarantee optimal results for routing. This aligns with observations from prior work [22], which found that well-calibrated confidence scores do not necessarily imply a strong correlation with predictions correctness. Overall, Self-REF achieves superior routing performance across all datasets and local LLMs, even though it only outperforms baseline methods under most of the calibration assessments.
MMLU OpenbookQA GSM8K MedQA Route Rate ECE BS CE Route Rate ECE BS CE Route Rate ECE BS CE Route Rate ECE BS CE Verbalizing Uncertainty 100% 0.217 0.333 7.383 75% 0.148 0.311 7.688 92% 0.047 0.307 7.999 100% 0.069 0.331 1.206 Verbalizing Yes/No Token 100% 0.466 0.397 10.281 95% 0.291 0.415 5.319 100% 0.107 0.384 8.369 100% 0.090 0.470 11.276 Zero-shot Logits 100% 0.347 0.418 1.733* 90% 0.225 0.496 1.705 90% 0.027 0.292* 3.312 100% 0.108 0.736 3.767 Finetuned Logits 90% 0.081 0.206 0.602 70% 0.095 0.238 0.676 70% 0.055 0.256 0.930 80% 0.059 0.265 0.844 Self-REF - Llama3-8B-Instruct 39% 0.040 0.320* 11.534 49% 0.090 0.254* 1.118* 65% 0.019 0.400 3.088* 40% 0.066* 0.278* 1.085*
MMLU OpenbookQA GSM8K MedQA Route Rate ECE BS CE Route Rate ECE BS CE Route Rate ECE BS CE Route Rate ECE BS CE Verbalizing Uncertainty 100% 0.126 0.313 8.725 100% 0.221 0.434 10.846 100% 0.163 0.669 3.822 100% 0.087 0.564 2.480 Verbalizing Yes/No Token 100% 0.297 0.345 4.380 100% 0.252 0.394 5.863 100% 0.146 0.821 23.98 100% 0.090 0.579 17.14 Zero-shot Logits 95% 0.311* 0.353 1.643 100% 0.232 0.442 2.498 100% 0.046 0.398 1.957 100% 0.087 0.507* 2.409 Finetuned Logits 95% 0.173 0.263 0.751 50% 0.143 0.207* 0.627 100% 0.050* 0.323 1.236* 90% 0.023 0.226 0.633 Self-REF - Mistral-7B-Instruct 70% 0.369 0.293* 1.021* 40% 0.157* 0.193 0.869* 75% 0.068 0.273 0.968 70% 0.037* 0.226 0.653*
Table 2: Confidence token prediction on Llama3-8B-Instruct and Mistral-7B-Instruct. “Route Rate” denotes the proportion of queries that must be routed to achieve performance comparable to stronger LLMs. “Gray block” refers to the best routing rate, “bold format” is the best calibration scores, and “*” is the second best calibration scores.
Figure 4: Ablation studies on gradient masking among Llama3-8B-Instruct and Mistral-7B-Instruct demonstrate its benefits for improving the utility of downstream tasks.
@@ 5.4 Ablation Study on Gradient Masking
To understand the impact of gradient masking on downstream fine-tuning performance, we perform an ablation study of Self-REF with and without gradient masking on the OpenbookQA dataset. For both Llama3-8B-Instruct and Mistral-7B-Instruct, we observe that gradient masking improves downstream routing task accuracy over not having gradient masking (Figure 4). This aligns with intuition since the purpose of gradient masking is to prevent the LLM from learning to make wrong predictions when trained on samples with <UN> tokens.
@@ 5.5 Transferability of Confidence Token
We further explore the transferability of Self-REF across unseen datasets (i.e., not included in the training data). With fine-tuned LoRA weights on OpenbookQA, we test on MMLU datasets for Llama3-8B-Instruct and Mistral-7B-Instruct. The results show that Self-REF outperforms the baselines with good routing performance on the MMLU dataset even when using the transferred LoRA weights of the OpenbookQA dataset (Figure 5). To achieve parity with the 70B model, we observe that 75% and 70% of routing rate are required on MMLU datasets under Mistral-7B-Instruct and Llama3-8B-Instruct, respectively. Furthermore, compared to the results obtained by direct fine-tuning on MMLU (i.e., a 70% routing rate on Mistral-7B-Instruct and a 40% routing rate on Llama3-8B-Instruct shown in Figure 2), transferred Self-REF achieves competitive routing performance with only a slight degradation in routing rates. This demonstrates the potential of Self-REF to be trained for general, multi-task purposes due to its transferability.
Figure 5: Experiment results on the transferability of Self-REF. Local LLMs are finetuned on OpenbookQA and inference the results on MMLU dataset.
@@ 6 Discussion
Self-REF is a lightweight approach that requires only LoRA fine-tuning of 8B and 7B LLMs, yet it delivers comparable or even superior system-level routing performance compared to the non-finetuned 70B model. Note that a non-finetuned 70B model is used to mimic the common setting where practitioners may not have sufficient resources to fine-tune such a large model, but can instead fine-tune a smaller model on the task of interest. Overall, the Self-REF technique outperforms all baselines, with fine-tuned logits achieving the next best performance. In this section, we dive into the following questions: (1) Why can Self-REF after confidence-based routing achieve better performance than the non-finetuned 70B model? and (2) How can one navigate the tradeoffs in routing and rejection?
@@
For question (1), we believe that one possible reason behind this phenomenon is that fine-tuning with Self-REF does not degrade the model’s performance and effectively introduces correct knowledge, such as niche astronomy knowledge or complicated moral disputes, that the model does not inherently possess. Additionally, since the error distributions of two independent models are unlikely to overlap perfectly, the ability to route can serve as a form of ensembling of models. In the Appendix, we provide some case studies from MMLU that the powerful Llama3-70B-Instruct model answered incorrectly, but Self-REF with the smaller Llama-8B model answered confidently and correctly. As shown in Appendix Figures 6, the Llama3-70B-Instruct may lack specific knowledge to judge a question of complicated moral disputes without fine-tuning and injecting the knowledge directly. More case studies are provided in Appendix 5.5.
@@
For question (2), one must consider the parameters of the downstream setting. When navigating the accuracy vs. routing rate trade-off, one must consider the setting’s cost of routing. For example, if one is utilizing an LLM on a rover on Mars, then the latency and monetary cost of routing to a larger LLM on a server on Earth is enormous, and one may want to avoid routing at all costs. If one has a limited budget for API calls to a more powerful LLM, then one may also want to route sparingly. On the other hand, if monetary cost is not an issue, and latency is not critical, then one can have the entire trade-off curve available to them, and simply choose the point that obtains the highest accuracy. When navigating the rejection learning setting, one must consider the risk of an incorrect answer. For example, if one is using an LLM to inform medical treatment, then one may choose to abstain from answering more often. On the other hand if one is using LLM for creative endeavors and correctness is not vital, then one may choose to let the LLM make its predictions and rarely abstain. Overall, these settings are highly customizable and controllable depending on the parameters that one is faced with in reality.
Through uncertainty-based routing or rejection, Self-REF is designed to improve overall system performance while reducing cost. In our analysis, we primarily consider cost through the lens of GPU-hours and latency during inference, ignoring network signal transmission times between different models on different devices. While these could also factor into the cost-performance tradeoffs, signal quality across devices depends on the congestion and bandwidth of the network, and so it is difficult to reliably account for it in the analysis. Another potential risk of routing to larger LLMs is the possibility of data leakage during transmission across the network. Routing user queries to different LLMs can also come at a cost to privacy, especially when data is transmitted to third-party or unauthorized models for inference. Further research into securing routing paths and protocols for communication between devices on private cloud servers could help reduce this type of routing cost.
@@ 7 Conclusion
In this work, we introduce Self-REF, a lightweight fine-tuning strategy that incorporates confidence tokens, enabling LLMs to better express their confidence levels in a way that closely correlates with prediction correctness. These confidence scores are valuable for confidence-based routing and rejection learning, which, as we show, can improve system performance, efficiency, and output reliability.
Our findings suggest that integrating confidence tokens into LLMs is a promising step toward improving their reliability and safety, especially in real-world applications where errors carry significant costs. Future research could study explicit incorporation of a rejection choice into the Self-REF fine-tuning process, rather than relying solely on thresholding confidence token scores for rejection learning.
@@
Appendix A Limitations and Potential Risks
Our proposed Self-REF is designed for self-routing purposes to achieve better downstream task performance. We primarily focus on the cost of GPU hours and latency in different sizes of the LLMs during inference, where we ignore the signal transmission time between different models on different devices. The quality of signal transmission between each device highly depends on the environment and congestion control of the bandwidth. Therefore, it is hard and complicated to consider the transmission time in the routing task. One of the potential risks of routing to larger LLMs is the possibility of data leakage during transmission. Moreover, routing user queries to different LLMs can raise privacy concerns, especially when data is transmitted to third-party or unauthorized models for inference. Further research into securing routing paths and designing network protocols for communication between devices on private clouds and servers is highly recommended.
Appendix B Related Work
B.1 Uncertainty Quantification in LLMs

### v3

@@ 4.2 Experiment Settings
Figure 3: ROC curves of Llama3-8B-Instruct on MMLU (a) and OpenbookQA (b); and Mistral-7B-Instruct on MMLU (c) and OpenbookQA (d), for the rejection learning task. The true positive rate (recall) indicates the proportion of samples where “none of the above” is the ground truth that is correctly rejected. The false positive rate is the proportion of original samples that have been falsely rejected. Across all settings, Self-REF outperforms other baselines in learning to reject.
Confidence-based Rejection Learning Task
Implementation Details
We select Llama-3-8B-Instruct (Touvron et al., 2023) and Mistral-7B-v0.3-Instruct (Jiang et al., 2023) models as the pre-trained local base LLMs for Self-REF. The Llama3-70B-Instruct model (Touvron et al., 2023) is the larger, more powerful LLM to which unconfident queries are routed.
Parameter-efficient fine-tuning is performed using LoRA adapters with ranks 2, 4, and 8 following the observations from (Hayou et al., 2024). Under at most eight epochs of training, all other hyperparameters on model training are decided by grid search with the perplexity loss on validation sets. For more details about the hyper-parameters, see Appendix C.
5 Experiment Results
5.1 Routing Performance (RQ1)
Overall Accuracy
Confidence-based routing using Self-REF consistently achieves the best accuracy vs. routing rate trade-off for all four datasets (MMLU, OpenbookQA, GSM8K, MedQA) and both local LLMs (Llama3-8B-Instruct model and Mistral-7B-Instruct) (Figure 2). Using Self-REF on Llama3-8B-Instruct for MMLU, confidence-based routing of just 39% of queries can achieve comparable performance to routing to the Llama3-70B-Instruct model alone. Similar observations hold for the OpenbookQA, GSM8K, and MedQA datasets, where routing the Llama3-8B-Instruct model with rates 49%, 65%, and 40%, respectively, can yield comparable performance to Llama3-70B-Instruct alone. For Self-REF on Mistral-7B-Instruct, the minimum routing rates to achieve parity with Llama3-70B-Instruct are 70%, 50%, 75%, and 70% for MMLU, OpenBookQA, GSM8K, and MedQA, respectively. Among the baselines, the fine-tuned logits tend to perform second-best.
Efficiency Gains
Smaller local models often have lower latency (sec) during inference as well as lower monetary cost. Thus, routing only a fraction of queries to larger, more powerful models can significantly improve the efficiency of the overall system compared to relying entirely on the more powerful models. For each small LLM (Llama3-8B-Instruct and Mistral-7B-Instruct), we select the best routing rate as the minimum routing rate that achieves comparable performance to Llama3-70B-Instruct (see Figure 2 for the trade-off curves). Without sacrificing performance compared to Llama3-70B-Instruct, we observe that Self-REF achieves as much as a 2.03×2.03$\times$2.03 × latency improvement for Llama3-8B-Instruct on MMLU and a 2.00×2.$0$ latency improvement for Mistral-7B-Instruct on OpenbookQA (as shown in Table 1). This indicates the efficiency advantage of leveraging Self-REF in routing tasks.
Table 2: Calibration metrics and routing rates of Self-REF and baselines using Llama3-8B-Instruct and Mistral-7B-Instruct, on the MMLU, OpenbookQA, GSM8K, and MedQA datasets. “Route” denotes the routing rate or proportion of queries that must be routed to achieve performance comparable to Llama3-70B-Instruct. The gray block refers to the best routing rate, “bold format” is the best calibration score, and “*” is the second best calibration score. Self-REF consistently achieves the best routing rate, but does not always achieve the best calibration score
Llama3-8B-Instruct MMLU OpenbookQA GSM8K MedQA Route ECE BS CE Route ECE BS CE Route ECE BS CE Route ECE BS CE Verbalizing Uncertainty 100% 0.217 0.333 7.383 75% 0.148 0.311 7.688 92% 0.047 0.307 7.999 100% 0.069 0.331 1.206 Verbalizing Yes/No Token 100% 0.466 0.397 10.281 95% 0.291 0.415 5.319 100% 0.107 0.384 8.369 100% 0.090 0.470 11.276 Zero-shot Logits 100% 0.347 0.418 1.733* 90% 0.225 0.496 1.705 90% 0.027 0.292* 3.312 100% 0.108 0.736 3.767 Finetuned Logits 90% 0.081 0.206 0.602 70% 0.095 0.238 0.676 70% 0.055 0.256 0.930 80% 0.059 0.265 0.844 Self-REF - Llama3-8B-Instruct 39% 0.040 0.320* 11.534 49% 0.090 0.254* 1.118* 65% 0.019 0.400 3.088* 40% 0.066* 0.278* 1.085*
Mistral-7B-Instruct MMLU OpenbookQA GSM8K MedQA Route ECE BS CE Route ECE BS CE Route ECE BS CE Route ECE BS CE Verbalizing Uncertainty 100% 0.126 0.313 8.725 100% 0.221 0.434 10.846 100% 0.163 0.669 3.822 100% 0.087 0.564 2.480 Verbalizing Yes/No Token 100% 0.297 0.345 4.380 100% 0.252 0.394 5.863 100% 0.146 0.821 23.98 100% 0.090 0.579 17.14 Zero-shot Logits 95% 0.311* 0.353 1.643 100% 0.232 0.442 2.498 100% 0.046 0.398 1.957 100% 0.087 0.507* 2.409 Finetuned Logits 95% 0.173 0.263 0.751 50% 0.143 0.207* 0.627 100% 0.050* 0.323 1.236* 90% 0.023 0.226 0.633 Self-REF - Mistral-7B-Instruct 70% 0.369 0.293* 1.021* 40% 0.157* 0.193 0.869* 75% 0.068 0.273 0.968 70% 0.037* 0.226 0.653*
@@ 5.2 Rejection Learning Performance (RQ2)
In this section, we analyze the utility of confidence scores from Self-REF for the purpose of learning to abstain from answering. Without incorporating additional knowledge of the rejection task during training, Self-REF and baselines rely solely on thresholding the confidence scores to determine whether to reject a prediction. In safety-critical applications where one may hope to detect cases that should be abstained from answering (i.e., high recall on the reject option) while retaining performance on the task at hand (i.e., low false positive rate), it is useful to examine the ROC curve, which makes this trade-off. Across all thresholds, we observe that on both MMLU and OpenbookQA dataset, Self-REF outperforms all baselines on the rejection learning task (Figure 3). Because the confidence tokens are trained and conditioned on the correctness of the predicted answers, Self-REF has a higher likelihood of being aware of the correctness of its generated responses using these tokens. In this manner, Self-REF may better distinguish when there is no correct answer to choose, resulting in better rejection learning capabilities.
@@ 5.3 Calibration with Utility (RQ3)
Here, we analyze the correlation between calibration, or how well predicted probabilities align with actual probabilities, and utility for downstream routing. Following the settings from Tian et al. (2023); Błasiok et al. (2023), three calibration metrics are used to assess the calibration: (1) expected calibration error, ECE (Guo et al., 2017); (2) Brier Score, BS (Brier, 1950); and (3) the cross-entropy score, CE. Alongside these calibration metrics, we also show the least rates of routing required for each technique to attain comparable performance to Llama3-70B-Instruct. As shown in Table 2, although Self-REF often achieves the best or second-best calibration scores, better-calibrated methods do not guarantee optimal results for routing. This aligns with observations from prior work (Huang et al., 2023), which found that well-calibrated confidence scores do not necessarily imply a strong correlation with the correctness of the predictions. Overall, Self-REF achieves superior routing performance across all datasets and local LLMs, even though it only outperforms baseline methods under most of the calibration assessments.
 Refer to caption     Refer to caption
Figure 4: Impact of gradient masking during Self-REF fine-tuning versus not (No Mask), assessed on the routing task.
 Refer to caption     Refer to caption
Figure 5: Transferability of Self-REF, assessed on the routing task. Local LLMs are fine-tuned with Self-REF on OpenbookQA and inference is run on the MMLU dataset.
Question-1: What is the difference between a male and a female catheter? Subject: Clinical Knowledge Choices: [A] Male and female catheters are different colors. [B] Male catheters are longer than female catheters. [C] Male catheters are bigger than female catheters. [D] Female catheters are longer than male catheters Ground Truth Answer: [B] Male catheters are longer than female catheters. Self-REF-8B: (✕) [D] Female catheters are longer than male catheters. <UN> $\rightarrow$ Route Larger LLM-70B: (✓) [B] Male catheters are longer than female catheters.
Question-2: A certain pipelined RISC machine has 8 general-purpose registers R0, R1, . . . , R7 and supports the following operations. [Truncated 50 words]. If the contents of these three registers must not be modified, what is the minimum number of clock cycles required for an operation sequence that computes the value of AB + ABC + BC? Subject: College Computer Science Choices: [A] 5 [B] 6 [C] 7 [D] 8 Ground Truth Answer: [B] 6. Self-REF-8B: (✕) [C] 7. <CN> $\rightarrow$ Overconfident Larger LLM-70B: (✓) [B] 6.
Question-3: This question refers to the following information. Albeit the king’s Majesty justly and rightfully is and ought to be the supreme head of the Church of England, [Truncated 230 words]; From the passage, one may infer that the English Parliament wished to argue that the Act of Supremacy would Subject: High School European History Choices: [A] give the English king a new position of authority [B] give the position of head of the Church of England to Henry VIII alone and exclude his heirs [C] establish Calvinism as the one true theology in England [D] end various forms of corruption plaguing the Church in England Ground Truth Answer: [D] end various forms of corruption plaguing the Church in England Self-REF-8B: (✓) [D] end various forms of corruption plaguing the Church in England. <UN> $\rightarrow$ Underconfident Larger LLM-70B: (✓) [D] end various forms of corruption plaguing the Church in England.
Figure 6: Examples of routing from Llama-8B-Instruct to Llama3-70B-Instruct, on the MMLU dataset.
@@ 5.4 Ablation Study on Gradient Masking
To evaluate the impact of gradient masking on downstream fine-tuning performance, we consider Self-REF with and without gradient masking on the OpenbookQA dataset. For both Llama3-8B-Instruct and Mistral-7B-Instruct, we find that applying gradient masking consistently improves routing task accuracy (Figure 4). This supports the intuition behind gradient masking: by masking gradients on samples with <UN> tokens, we prevent the model from learning spurious patterns based on uncertain inputs.
@@ 5.5 Transferability of Confidence Token
We further explore the transferability of Self-REF across unseen datasets (i.e., not included in the training data). With fine-tuned LoRA weights on OpenbookQA, we test on MMLU for Llama3-8B-Instruct and Mistral-7B-Instruct. Self-REF outperforms the baselines with good routing performance on the MMLU dataset even when using the transferred LoRA weights of the OpenbookQA dataset (Figure 5). To achieve parity with the 70B model, we observe that routing rates of 75% and 70% are required on MMLU under Mistral-7B-Instruct and Llama3-8B-Instruct, respectively. Furthermore, compared to the direct fine-tuning on MMLU (i.e., a 70% routing rate on Mistral-7B-Instruct and a 40% routing rate on Llama3-8B-Instruct, Figure 2), transferred Self-REF achieves competitive routing performance with only a slight degradation in routing rates. This demonstrates the potential of Self-REF to be trained for general, multi-task purposes due to its transferability.
5.6 Case Studies
While Self-REF significantly outperforms other methods in routing accuracy, some failure cases still occur, particularly on niche subjects or tasks requiring nuanced reasoning. In underconfident cases, the model often hesitates on context-heavy (i.e., High School Europe History in Figure 6) or ambiguous subjects, highlighting its sensitivity to topics requiring broader reasoning. Overconfident failure cases occur in technical or highly structured domains (i.e., College Computer Science in Figure 6), where the model appears familiar but fails to grasp nuanced or edge-case reasoning. This phenomenon is also observed in the jurisprudence domain, where subtle distinctions in legal language and context-specific reasoning often lead to overconfident yet incorrect predictions. Conversely, correctly predicted confident cases typically involve broad conceptual topics, due to training coverage and clearer correctness signals. Meanwhile, incorrect but uncertain cases often correspond to detail-intensive subjects, where the model appropriately expresses hesitation due to limited recall or understanding. For more details see Appendix E.
@@ 6 Discussion
Self-REF delivers comparable or even superior system-level routing performance compared to the non-finetuned 70B model. Note that a non-finetuned 70B model is used to mimic the common setting where practitioners may not have sufficient resources to fine-tune such a large model, but can instead fine-tune a smaller model on the task of interest. In this section, we dive into the following questions: (1) Why can Self-REF after confidence-based routing achieve better performance than the non-finetuned 70B model? and (2) How can one navigate the tradeoffs in routing and rejection?
@@
For question (1), we believe that one possible reason behind this phenomenon is that fine-tuning with Self-REF does not degrade the model’s performance and effectively introduces correct knowledge, such as niche astronomy knowledge or complicated moral disputes, that the model does not inherently possess. Additionally, since the error distributions of two independent models are unlikely to overlap perfectly, the ability to route can serve as a form of ensembling of models. In Appendix D, we provide some case studies from MMLU that the powerful Llama3-70B-Instruct model answered incorrectly, but Self-REF with the smaller Llama-8B model answered confidently and correctly. As shown in Appendix Figure 8, the Llama3-70B-Instruct may lack specific knowledge to judge a question of complicated moral disputes without fine-tuning and injecting the knowledge directly. More case studies are provided in Appendix E.
For question (2), Self-REF provides a continuous-valued confidence score which enables downstream decision-makers a granular level of control. By contrast, we observe that certain methods for extracting confidence scores, such as verbalizing uncertainty, can suffer from mode collapse, where the scores cluster around a limited set of numbers (Figure 2, red dotted triangle lines). To appropriately leverage the granular control provided by Self-REF, one must consider the parameters of the downstream setting.
@@
When navigating the accuracy vs. routing rate trade-off, one must consider the setting’s cost of routing. For example, if one is utilizing an LLM on a rover on Mars, then the latency and monetary cost of routing to a larger LLM on a server on Earth is enormous, and one may want to avoid routing at all costs. If one has a limited budget for API calls to a more powerful LLM, then one may also want to route sparingly. On the other hand, if monetary cost is not an issue, and latency is not critical, then one can have the entire trade-off curve available to them, and simply choose the point that obtains the highest accuracy.
When navigating the rejection learning setting, one must consider the risk of an incorrect answer. For example, if one is using an LLM to inform medical treatment, then one may choose to abstain from answering more often. On the other hand if one is using LLM for creative endeavors and correctness is not vital, then one may choose to let the LLM make its predictions and rarely abstain. Overall, these settings are highly customizable and controllable depending on the parameters that one is faced with in reality.
Through confidence-based routing and rejection, Self-REF enhances the overall performance and safety of a system while reducing cost. In our analysis, we measure cost primarily in terms of GPU hours and latency, excluding network transmission time between models which could be hosted on different machines (which is subject to varying signal quality, setup costs, and network congestion). Future research in downstream applications could seek to better characterize these networking costs, which may contribute substantially to savings gained from confidence-based routing.
@@ 7 Conclusion
In this work, we introduce Self-REF, a lightweight fine-tuning strategy that incorporates confidence tokens, enabling LLMs to better express their confidence levels in a way that closely correlates with prediction correctness. These confidence scores are valuable for confidence-based routing and rejection learning, which can improve system performance, efficiency, and output reliability for both LLM routing and LLM rejection learning purposes. Our findings suggest that integrating confidence tokens into LLMs is a promising step toward improving their reliability and safety, particularly in real-world applications where errors carry significant costs.
More broadly, Self-REF opens new avenues for research into confidence-based routing and rejection. For instance, Self-REF could be extended to route queries to multiple specialized LLMs rather than a single powerful model. Another direction involves finer-grained annotation schemes, where individual sentences are labeled as confident or uncertain. Based on confidence, the system could route to different models, stop early, or re-invoke the reasoning processes to refine outputs. For confidence-based rejection, one could expand Self-REF to incorporate explanations of rejections or to suggest actionable clarifying information.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@
Appendix A Related Work
A.1 Uncertainty Quantification in LLMs